bdrammeh:

A circular loop of radius 0.0400 m is oriented 75.0 degrees to a magnetic field. If the magnetic flux through the loop is 9.59 x 10^-7 Wb, what is the strength of the field? ___x10^__-T

imqwerty:

any ideas?

bdrammeh:

of wat

imqwerty:

do you know how to calculate magnetic flux

bdrammeh:

no

imqwerty:

it's essentially just the strength of magnetic field multiplied by the area it's acting upon

imqwerty:

since your circular loop is tilted at an \(75^o\) to the direction of the magnetic field you can't just multiply the area of the circle with the field strength to get the magnetic flux

bdrammeh:

can u do it

imqwerty:

you're gonna do the calculations

bdrammeh:

but can u plug it in

imqwerty:

you'll have to take the component of the area which is perpendicular to the direction of the field and multiply it with field strength to get magnetic flux

bdrammeh:

just do it

imqwerty:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bdrammeh just do it \(\color{#0cbb34}{\text{End of Quote}}\) k

imqwerty:

\(Area \times \sin(angle~with~the~field) \times strength~of~field = magnetic~flux\)

bdrammeh:

plug in the number s thats all

imqwerty:

why don;t you do it? isn't it your homework?

bdrammeh:

ya but just help me out real quick

imqwerty:

are you not able to plug in the values

bdrammeh:

i did plug it the values but threa nswer still seems to be wrong

TheSmartOne:

What is the area of a circular loop of radius 0.04 m? Multiply that by sin 75 And then magnetic flux divided by that number that you got previously will give you the strength of the field Also if you show us your work, it'll be easier for us to see where you're making the mistake and help you better :)

bdrammeh:

i dont the the magnetic flux and dont know the answer

TheSmartOne:

How do you find the area of a circle? Do you know the formula? They gave the magnetic flux in the question itself. Reread the question.

TheSmartOne:

If you work with us, we can solve this question much much faster. But you have to tell us what you know, where you're stuck, and how far you've gotten

bdrammeh:

just tell me the step by step for this question and i can solve the rest or you just acn send me the link of the calculator to use

TheSmartOne:

I just broke it down step by step I'm not sure what's making this question difficult for you You have a formula You are given all these values and you plug them in You have one variable left which you solve for Do you understand that? Now, do you know the formula for finding the area of a circle? Can you find the area if r = 0.04 m Multiply that by sin(75) Now whatever number you get, let's call that 'x' Idk this number rn, you have to find it Now you do magnetic flux/x Magnetic flux is given in the question as 9.59*10^-7 And then you're done

TheSmartOne:

Wait apparently the equation is magnetic flux = magnetic field * area * cos \(\theta\) So you should be doing cos(75)